OpenStudy (anonymous):

y=x^2-x, x>=1/2 solve for f^-1(x) (inverse)

OpenStudy (anonymous):

y=x^2-x+1/4-1/4 -->y=(x-1/2)^2-1/4 ---> x=+-(y+1/4)+1/2 -->(inverse) y= +-(x+1/4)+1/2

OpenStudy (anonymous):

so you added 1/4 and -1/4 but how did you go from step 1 to step 2

OpenStudy (anonymous):

I had done it because I wanted to have the parameter x alone in one side 'n' others another side. Ok?

OpenStudy (anonymous):

y=x^2-x+1/4-1/4 -->y=(x-1/2)^2-1/4 what happened inbetween?

OpenStudy (anonymous):

(x-1/2)^2-1/4=(x^2-x+1/4) -1/4=x^2-x

OpenStudy (anonymous):

Ok, I have to go!